OpenStudy (anonymous):

If y(x) = x/5, what is x when y(x) = 5 ? a) 25 b) 1 c) 0 d) 10

OpenStudy (anonymous):

a

OpenStudy (anonymous):

25

OpenStudy (anonymous):

If y = x/5 and y = 5 Then x/5 = 5